# Equilibrium in Style: A Modeling Framework on the Cash Flow and the Life Cycle of a Consumer Store

### v1

@@ 3.2 Style supply
Next, we will discuss the strategy and the supply of store owners. In our model, the store owner opens at time $t$ and needs to decide on the style ${\boldsymbol{z}}$ and the price of the produced products $\theta,$ which cannot be changed afterward. Kashyap (1995) shows that the change in retail prices is very slow. Bils and Klenow (2004) provides evidence that the prices of services like haircuts for males and beauty parlors for females change only every 20 months or more. So we first assume that price and style cannot be changed after opening for most of the stores. In Section 4.4 we will consider situations where the style and price of particular stores can be changed over time. Moreover, storefront attributes are related to the initial investment and we consider them as exogenous variables. The store needs to satisfy constraint $g_{I}(\boldsymbol{\xi})\leqslant I,$ where $I$ is a fixed initial investment for a given store owner. $g_{I}(\cdot)$ is a smooth convex cost function. Denote that $G_{I}(\boldsymbol{z})=G_{I}\left(\boldsymbol{x},\boldsymbol{\xi}\right)=\boldsymbol{x}^{\prime}\cdot\boldsymbol{0}+g_{I}(\boldsymbol{\xi})$ and the constraint is also denoted as $G_{I}(\boldsymbol{z})\leqslant I$. Furthermore, we simply assume that this constraint is tight for all types of consumers, i.e. $g_{I}\left(\hat{\boldsymbol{\xi}}_{j,t}\right)>I,$ for $j=1,2,\ldots,m,$ $t>0$ and $I\in\mathcal{I}.$ Let the set of all possible initial investments be $\mathcal{I}=[I_{m},I_{M}]\subset\mathbb{R}^{+},$ which is a bounded set of a non-zero measure that does not change over time. To maximize the cash flow at time $t,$ the store owner’s optimization goal is written as,
@@
is the set of non-isolated points of the probability density. This implies that only styles that form a scale will lead to effective purchasing behavior among consumers, otherwise the purchase probability corresponding to one or several isolated styles is $0$ When the density is non-zero on the full space, for example $\rho_{jt}\in\mathscr{P}_{jt}\ (j=1,2,\ldots,m),$ then $C_{t}=\mathbb{R}^{p+q}\times\mathbb{R}^{+}$ and the second constraint can be removed. The objective in (21) is called the cash flow density. If the solution of (21) exists and is unique, we denote it as $\left(\boldsymbol{z}_{t}^{(s)}(I),\theta\left(\boldsymbol{z}_{t}^{(s)}(I)\right)\right).$ When the cash flow density drops to a certain threshold $r(I),$ the store will shut down. So given $I,$ we can define the longest lifespan of a store that exists at time $t$ as
@@ 3.3 Optimal strategy uniqueness for supply
(21) can be a non-convex optimization problem, so the solution to this optimization equation is not necessarily unique and even one style may correspond to multiple optimal prices. It is natural to assume that for a given style, only the infimum of the optimal prices will occur, otherwise, consumers would be perfectly capable of choosing a lower-priced store of the same style. Thus, solving (21) we obtain the set of optimal styles
@@
of the store that opens at time $t$ with initial investment $I,$ where $\Lambda_{t}(I)$ is the set of indicators. This can be considered the supply of the store owner.
@@
Consider an illustrative example in the restaurant industry, where there are young and old people in the economy, $\frac{1}{3}$ and $\frac{2}{3}$ of the economy, respectively. At time $0,$ the young like $3$g of salt in their food best, while the old like $1$g of salt best, and they both like it best when there are $5$ lights at the entrance of the restaurant, which costs \$$10$ each. Their purchase probability densities are
@@
For some extreme density (even if bounded), (21) may be unsolvable, i.e. it is not possible to define the style that the store chooses. However, when $\rho_{jt}({\boldsymbol{z}}|\theta)\in\mathscr{P}_{jt},$ the following proposition guarantees the existence of the solution.
@@
Due to the presence of different types of consumers, even stores open in the same time and with the same initial investment are likely to be dispersed in terms of style and price, which is consistent with the reality. However, there is generally a mainstream, usually designed to target sales volume to certain specific types of consumers. To capture this, and also for simplicity, we assume that the store owner makes decisions based on the importance of the purchase probability for each type of consumer.
@@
The implication of assumption 2 is that with homogeneou cash flow, the store will make decisions based on the purchase probability of each type of consumers. From the first to the $m$th type of consumers, their purchase probability is of decreasing importance to the store. That is, stores are more inclined to cater to the preference of the first few types of consumers, thereby greatly maximizing sales volume on them, rather than the latter types. Proposition 3.3 shows that ${\boldsymbol{z}}_{t}^{(s)}(I)$ is unique for a broad class of purchase probability density forms that we concerned.
@@
If there exists $j_{0}\in\{1,2\ldots,m\}$ s.t. given $\theta$ the maximum point of $\rho_{j_{0}t}({\boldsymbol{z}}|\theta)$ on $D=\left\{{\boldsymbol{z}}\in\mathbb{R}^{p+q}:G_{I}(\boldsymbol{z})\leqslant I\right\}$ is unique, then #$\#(S_{m})=1.$
@@
is equivalent to (36), of which the solution is unique, see Boyd and Vandenberghe (2004). ∎
@@ 4.1 Preference shifting
Simonson and Tversky (1992) illustrates that consumers’ preferences for products are influenced by information, advertising, and other market factors. Although the shift in product preferences may be non-linear, it is slower compared to the life cycle of stores. Our assumption uses a first-order approximation of the shift. The following proposition states that under this Assumption 3, the scoring function for each type of consumers has time translation invariance.
@@
Let $\rho_{jt}({\boldsymbol{z}}|\theta)=K_{jt}\phi\left(C_{j}-\frac{\lambda_{j}}{2}\left\|\boldsymbol{z}-\hat{\boldsymbol{z}}_{j,t}\right\|^{2}\right)\cdot e^{-\gamma_{j}\theta}.$ There exists bounded positive function $\tilde{B}_{t}(\cdot)\leqslant B_{t}(\cdot)$ such that on
@@
for $j=2,3,\ldots,m,$ $t^{\prime}\in(t+t_{\varepsilon},\infty)$ for some $t_{\varepsilon}>0$ and $I\in\mathcal{I}$
@@
We still consider illustrative example in Example 3.1. But now half of the old people are younger (called middle-aged) with heavier tastes, and they prefer $5$g of salt. We envision that the tastes of these three types of consumers change over time, and that their most preferred amount of salt increased by $1$g per year. Their scoring functions at time $t$ are
@@
Consider $K_{j0}\check{F}_{j}\left(x_{0}^{(s)}(20),\xi_{0}^{(s)}(20),\theta\left(\boldsymbol{z}_{0}^{(s)}(20)\right)\right)$ as an approximate long-run cumulative value of the purchase probability density. When $K_{10}$ is relatively large compared to $K_{20}$ and $K_{30}$, the first type of consumers become the dominant consumers, and we are going to look at the interaction between the other two types of consumers in this case. Figure 5 shows the case where we fix $K_{10}=1,$ and $K_{20}$ and $K_{30}$ vary in a range less than $1$, respectively. We can see that the purchase probability levels of the two have little cross-influence on each other.
@@
For the story told in Propositions 3.5 and 4.2 to happen, it is not in fact necessary that the level of the purchase probability of the dominant consumers is really high. As Proposition 4.3 show, when a type of consumers are a large proportion of the population, even a low level of purchase probability can make stores cater to their preference.
@@ 4.2 Equilibrium under preference shifting
From a game perspective, there are two parts of players in the game. There are infinitely many (uncountable) store owners denoted as $\mathcal{T}\times\mathcal{I},$ and for consumers, since the decisions at time $t_{1}$ and $t_{2}\ (t_{1}\neq t_{2})$ are independent (maximizing the utility at time $t_{1}$ and $t_{2},$ respectively), consumers are also set to have infinitely many (uncountable) denoted as $\mathcal{T}\times\{1,2,\ldots,m\}.$ The action set of the store owner $(t,I)$ is $\left\{(\boldsymbol{z},\theta)\in\mathbb{R}^{p+q}\times\mathbb{R}^{+}:G_{I}(\boldsymbol{z})\leqslant I\right\}$, and the action set of the consumer $(t,j)$ is $\left\{\rho:\int_{\mathbb{Z}_{t}}\rho\mathrm{d}\boldsymbol{z}=1\right\}.$ The payoff for the store owner $(t,I)$ is $\theta\sum_{j=1}^{m}\mathbf{P}^{(j)}\rho_{jt}({\boldsymbol{z}}|\theta)\cdot\mathds{1}_{\{(\boldsymbol{z},\theta)\in C_{t}\}},$ and the payoff for the consumer $(t,j)$ is $U_{jt}(\rho).$
This game has an uncountable number of players and the action set of consumers at time $t$ depends on the strategies of the store owners in the history prior to $t.$ Also, a store owner’s optimal strategy exist but is discontinuous with respect to the consumers’ action (see Example 3.1). These lead us not to assert the existence of an equilibrium. However, Theorem 4.1 provides a Nash equilibrium (Nash (1951)) for this model.
We denote the style, price, style set, and consumer purchase probability density under equilibrium as ${\boldsymbol{z}}_{t}^{*}(I),\theta({\boldsymbol{z}}_{t}^{*}(I)),\mathbb{Z}^{*}_{t},\rho^{*}_{jt},$ respectively.
@@
We know that $\boldsymbol{z}^{*}_{t}(I)=\boldsymbol{z}^{*}_{0}(I)+t\boldsymbol{d},$ and $\theta\left(\boldsymbol{z}^{*}_{t}(I)\right)=\theta\left(\boldsymbol{z}^{*}_{0}(I)\right).$ After some algebra, we get (68) and (70.)
Step2. For type $j$th consumers, the optimal purchase probability density satisfies
@@
Even if there is cooperation between stores opening in the same time, the equilibrium in Theorem 4.1 holds.
From Theorem 4.1 we can see some characteristics of the equilibrium state. First, the styles produced by the store also change in the direction of preference drift, while prices are only related to the initial investment of the store. Second, the purchase probability density for each type of consumers can be expressed as a monotonic decreasing function of the distance of the style from the optimal preference style, which indicates that the probability of a consumer purchasing from a fixed store decreases over time. Finally, the lifetime of the store is only related to its initial investment and not to the time when it opens.
@@ 4.3 Exponential transform function and theoretical conversion rate in equilibrium
$\tilde{\beta}_{jt}({\boldsymbol{z}}^{*}_{0}(I))$ $=\tilde{\beta}_{j0}({\boldsymbol{z}}^{*}_{0}(I))\frac{\int_{\Delta_{0}(\boldsymbol{z}_{0}(I),\varepsilon)}\rho^{*}_{jt}({\boldsymbol{z}}_{1})\mathrm{d}{\boldsymbol{z}}_{1}}{\int_{\Delta_{t}(\boldsymbol{z}_{0}(I),\varepsilon)}\rho^{*}_{j0}({\boldsymbol{z}}_{1})\mathrm{d}{\boldsymbol{z}}_{1}}.$ (89)
@@
When $t$ is sufficiently large, $\tilde{\beta}_{t}$ monotonically decreases with respect to $t$ and approaches 00. This is mainly due to the product style being eliminated by the market due to the style-shifting. However, due to the dispersion of preferences among different types of people, $\tilde{\beta}_{t}$ may experience some fluctuations before decrease. Figure 6 shows the conversion rate curves of several different shapes. The following theorem demonstrates that the $\tilde{\beta}_{t}$ undergoes at most one growth before decreasing to 00 when the product preferences of different types are not dispersed. That is to say, only (1) and (2) in Figure 6 will appear when people’s product preferences are sufficiently concentrated. In other words, the $\tilde{\beta}_{t}$ curve has the shape of a rainbow or is a part of it.
@@ 4.4 Style updating
Let us consider that in the equilibrium in Theorem 4.1 a store which opens at time $t$ with initial investment $I$ has the ability to update style. There is a cost to update style, and the store budgets $S>0$ for the cost of style updating. From time 00 to time $t>0$, the store can change its style from $\boldsymbol{z}^{*}_{0}(I)$ to $\boldsymbol{z}^{*}_{0\rightarrow t}(I)$ smoothly. However, it needs to satisfy constraint

### v2

@@ 3.2 Style supply
Next, we will discuss the strategy and the supply of store owners. In our model, the store owner opens at time $t$ and needs to decide on the style ${\boldsymbol{z}}$ and the price $\theta,$ which cannot be changed afterward. Kashyap (1995) shows that the change in retail prices is very slow. Bils and Klenow (2004) provides evidence that the prices of services like haircuts for males and beauty parlors for females change only every 20 months or more. So we first assume that price and style cannot be changed after opening for most of the stores. In Section 4.4 we will consider situations where the style and price of particular stores can be changed over time. Moreover, storefront attributes are related to the initial investment and we consider the initial investment as exogenous variables. The store owner needs to satisfy constraint $g_{I}(\boldsymbol{\xi})\leqslant I,$ where $I$ is a fixed initial investment for a given store owner. $g_{I}(\cdot)$ is a smooth convex cost function. Denote that $G_{I}(\boldsymbol{z})=G_{I}\left(\boldsymbol{x},\boldsymbol{\xi}\right)=\boldsymbol{x}^{\prime}\cdot\boldsymbol{0}+g_{I}(\boldsymbol{\xi})$ and the constraint is also denoted as $G_{I}(\boldsymbol{z})\leqslant I$. Furthermore, we simply assume that this constraint is tight for all types of consumers, i.e. $g_{I}\left(\hat{\boldsymbol{\xi}}_{j,t}\right)>I,$ for $j=1,2,\ldots,m,$ $t>0$ and $I\in\mathcal{I}.$ To maximize the cash flow at time $t,$ the store owner’s optimization goal is written as,
@@
is the set of non-isolated points of the probability density domain. This implies that only styles that form a scale will lead to effective purchasing behavior among consumers, otherwise the purchase probability corresponding to one or several isolated styles is $0$ When the density is non-zero on the full space, for example $\rho_{jt}\in\mathscr{P}_{jt}\ (j=1,2,\ldots,m),$ then $C_{t}=\mathbb{R}^{p+q}\times\mathbb{R}^{+}$ and the second constraint can be removed. We call the objective in Problem (21) the cash flow density. If the solution of Problem (21) exists and is unique, we denote it as $\left(\boldsymbol{z}_{t}^{(s)}(I),\theta\left(\boldsymbol{z}_{t}^{(s)}(I)\right)\right).$ When the cash flow density drops to a certain threshold $r(I),$ the store will shut down. Given $I,$ we define the longest lifespan of a store that exists at time $t$ as
@@ 3.3 Optimal strategy uniqueness for supply
Problem (21) can be a non-convex optimization problem, so the solution to this optimization equation is not necessarily unique and even one style may correspond to multiple optimal prices. It is natural to assume that for a given style, only the infimum of the optimal prices will occur, otherwise, consumers would be perfectly capable of choosing a lower-priced store of the same style. Thus, in solving Problem (21), we obtain the set of optimal styles
@@
of the store that opens at time $t$ with initial investment $I,$ where $\Lambda_{t}(I)$ is the set of indicators. These can be considered the supply of the store owner.
@@
Consider an illustrative example in the restaurant industry, where there are young and old people in an economy, $\frac{1}{3}$ and $\frac{2}{3}$ of the population, respectively. At time $0,$ the young like $3$g of salt in their food best, while the old like $1$g of salt best, and they both like it best when there are $5$ lights at the entrance of the restaurant, which costs \$$10$ each. Naturally, their generalized purchase probability densities are
@@
For some extreme density (even if bounded), Problem (21) may be unsolvable, i.e. it is not possible to define the style that the store chooses. However, when $\rho_{jt}({\boldsymbol{z}}|\theta)\in\mathscr{P}_{jt},$ the following proposition guarantees the existence of the solution.
@@
Due to the presence of different types of consumers, even stores open in the same time and with the same initial investment are likely to be dispersed in terms of style and price, which is consistent with the reality. However, among these styles there is generally a mainstream, usually designed to target sales volume to specific types of consumers. To capture this, and also for simplicity, we assume that the store owner makes decisions based on the importance of the purchase probability for each type of consumer.
@@
The implication of Assumption 2 is that with homogeneous cash flow, the store will make decisions based on the purchase probability of each type of consumers. From the first to the $m$-th type of consumers, their purchase probability is of decreasing importance to the store. That is, stores are more inclined to cater to the preference of the first few types of consumers, thereby greatly maximizing sales volume on them, rather than the latter types. Proposition 3.3 shows that ${\boldsymbol{z}}_{t}^{(s)}(I)$ is unique for a broad class of purchase probability density forms that we concerned.
@@
If there exists $j_{0}\in\{1,2\ldots,m\}$ s.t. given $\theta$ the maximum point of $\rho_{j_{0}t}({\boldsymbol{z}}|\theta)$ on $D=\left\{{\boldsymbol{z}}\in\mathbb{R}^{p+q}:G_{I}(\boldsymbol{z})\leqslant I\right\}$ is unique, then #$\#(S_{m})\leqslant 1.$
@@
is equivalent to (36), of which the solution is unique; see Boyd and Vandenberghe (2004). ∎
@@ 4.1 Preference shifting
Simonson and Tversky (1992) illustrates that consumers’ preferences for products are influenced by information, advertising, and other market factors. Although the shift in product preferences may be non-linear, it is slower compared to the lifespan of consumer stores. Our assumption uses a first-order approximation of the shift. The following proposition states that under this Assumption 3, the scoring function for each type of consumers has time translation invariance.
@@
Let $\rho_{jt}({\boldsymbol{z}}|\theta)=K_{jt}\phi\left(C_{j}-\frac{\lambda_{j}}{2}\left\|\boldsymbol{z}-\hat{\boldsymbol{z}}_{j,t}\right\|^{2}\right)\cdot e^{-\gamma_{j}\theta}\in\mathcal{P}_{jt}.$ There exists bounded positive function $\tilde{B}_{t}(\cdot)\leqslant B_{t}(\cdot)$ such that on
@@
for $j=2,3,\ldots,m,$ $t^{\prime}\in(t+t_{\varepsilon},\infty)$, some $t_{\varepsilon}>0$ and $I\in\mathcal{I}$
@@
We still consider illustrative example in Example 3.1. But now half of the old people are younger (called middle-aged) with heavier tastes, and they prefer $5$g of salt. We envision that the tastes of these three types of consumers change over time and that their most preferred amount of salt increased by $1$g per year. Their scoring functions at time $t$ are
@@
Consider $K_{j0}\check{F}_{j}\left(x_{0}^{(s)}(20),\xi_{0}^{(s)}(20),\theta\left(\boldsymbol{z}_{0}^{(s)}(20)\right)\right)$ as an approximate long-run cumulative value of the generalized purchase probability density. When $K_{10}$ is relatively large compared to $K_{20}$ and $K_{30}$, the first type of consumers become the dominant consumers, and we are going to look at the interaction between the other two types of consumers in this case. Figure 5 shows the case where we fix $K_{10}=1,$ and $K_{20}$ and $K_{30}$ vary in a range less than $1$, respectively. We can see that the purchase probability levels of the two have little cross-influence on each other.
@@
For the story told in Propositions 3.5 and 4.2 to happen, it is not in fact necessary that the level of the generalized purchase probability of the dominant consumers is really high. As Proposition 4.3 shows, when a type of consumers is a large proportion of the population, even a low level of generalized purchase probability can make stores cater to their preference.
@@ 4.2 Equilibrium under preference shifting
From the perspective of game theory, we characterize this economy as a complete information game. There are two parts of players in the game. There are infinitely many (uncountable) store owners denoted as $\mathcal{T}\times\mathcal{I}$. For consumers, since the decisions at time $t_{1}$ and $t_{2}\ (t_{1}\neq t_{2})$ are independent (maximizing the utility at time $t_{1}$ and $t_{2},$ respectively), consumers are also set to be infinitely many (uncountable) denoted as $\mathcal{T}\times\{1,2,\ldots,m\}.$ The action set of the store owner $(t,I)$ is $\left\{(\boldsymbol{z},\theta)\in\mathbb{R}^{p+q}\times\mathbb{R}^{+}:G_{I}(\boldsymbol{z})\leqslant I\right\}$, and the action set of the consumer $(t,j)$ is $\left\{\rho:\int_{\mathbb{Z}_{t}}\rho\mathrm{d}\boldsymbol{z}=1\right\}.$ The payoff for the store owner $(t,I)$ is $\theta\sum_{j=1}^{m}\mathbf{P}^{(j)}\rho_{jt}({\boldsymbol{z}}|\theta)\cdot\mathds{1}_{\{(\boldsymbol{z},\theta)\in C_{t}\}},$ and the payoff for the consumer $(t,j)$ is $U_{jt}(\rho).$
This game has an uncountable number of players and the action set of consumers at time $t$ depends on the strategies of the store owners in the history prior to $t.$ Also, a store owner’s optimal strategy exists but is discontinuous with respect to the consumers’ action (see Example 3.1). These lead us not to assert the existence of an equilibrium. However, Theorem 4.1 provides a Nash equilibrium (Nash (1951)) for this model.
We denote the style, price, provided style set, and consumer purchase probability density under equilibrium as ${\boldsymbol{z}}_{t}^{*}(I),\theta({\boldsymbol{z}}_{t}^{*}(I)),\mathbb{Z}^{*}_{t},\rho^{*}_{jt},$ respectively.
@@
We know that $\boldsymbol{z}^{*}_{t}(I)=\boldsymbol{z}^{*}_{0}(I)+t\boldsymbol{d},$ and $\theta\left(\boldsymbol{z}^{*}_{t}(I)\right)=\theta\left(\boldsymbol{z}^{*}_{0}(I)\right).$ After some algebra, we get (68) and (70).
Step2. For type $j$-th consumers, the optimal purchase probability density satisfies
@@
Even if there is cooperation between stores opening at the same time, the equilibrium in Theorem 4.1 holds.
From Theorem 4.1 we can see some characteristics of the equilibrium state. First, the styles provided by newly entered stores also shift linearly in the direction of preference shifting over time, while prices are only related to the initial investment of the stores. Second, the purchase probability density for each type of consumers can be expressed as a monotonic decreasing function of the distance of the style from the optimal preference style, which indicates that the probability of a consumer purchasing from a fixed store will decrease to zero over time. Finally, this equilibrium is time stable and the lifespan of a store is only related to its initial investment and not to the time when it opens.
@@ 4.3 Exponential transform function and theoretical conversion rate in equilibrium
$\tilde{\beta}_{jt}({\boldsymbol{z}}^{*}_{0}(I))$ $=\tilde{\beta}_{j0}({\boldsymbol{z}}^{*}_{0}(I))\frac{\int_{\Delta_{t}(\boldsymbol{z}_{0}(I),\varepsilon)}\rho^{*}_{jt}({\boldsymbol{z}}_{1})\mathrm{d}{\boldsymbol{z}}_{1}}{\int_{\Delta_{0}(\boldsymbol{z}_{0}(I),\varepsilon)}\rho^{*}_{j0}({\boldsymbol{z}}_{1})\mathrm{d}{\boldsymbol{z}}_{1}}.$ (89)
@@
When $t$ is sufficiently large, $\tilde{\beta}_{t}$ monotonically decreases with respect to $t$ and approaches 00. This is mainly due to the product style being eliminated by the market due to the style-shifting. However, due to the dispersion of preferences among different types of people, $\tilde{\beta}_{t}$ may experience some fluctuations before decreasing. Figure 6 shows the conversion rate curves of several different shapes. The following theorem demonstrates that the $\tilde{\beta}_{t}$ undergoes at most one growth before decreasing to 00 when the product preferences of different types are not dispersed. That is to say, only (1) and (2) in Figure 6 will appear when people’s product preferences are sufficiently concentrated. In other words, the $\tilde{\beta}_{t}$ curve has the shape of a rainbow or is a part of it.
@@ 4.4 Style updating
Let us consider that in the equilibrium in Theorem 4.1 a store which opens at time $t$ with initial investment $I$ has the ability to update the style. There is a cost to update style, and the store budgets $S>0$ for the cost of style updating. From time 00 to time $t>0$, the store can change its style from $\boldsymbol{z}^{*}_{0}(I)$ to $\boldsymbol{z}^{*}_{0\rightarrow t}(I)$ smoothly. However, it needs to satisfy constraint